which one of the following has the lowest viscosity


A. Water

B. Motor oil

C. Marble syrup

D. Glue

E. None of the above


1
Expert's answer
2022-04-19T12:06:11-0400

A. Water: "0.001\space Pa\cdot s"

B. Motor oil: "0.005 - 4\space Pa\cdot s"

C. Marble syrup: "0.03-0.7 \space Pa\cdot s"

D. Glue (PVA): "10-18\space Pa\cdot s"


Answer. A. Water.